Determine whether the equation is an identity, a conditional equation or an inconsistent equation.

-7x+2/3 + 4/3 = -7x/2

Answer by stanbon(75887)

Determine whether the equation is an identity, a conditional equation or an inconsistent equation.
-7x+(2/3) + 4/3 = (-7x)/2
-7x + 2 = (-7x/2)
Multiply thru by 2 to get:
-14x + 4 = -7x
Add 14x to both sides to get:
7x = 4
x = 4/7
================
Cheers,
Stan H.